OpenStudy (anonymous):

88-18=46/9(h-6). solve for h

OpenStudy (anonymous):

70 = 46/9(h-6) h-6 =46/9*70 h = 46/9*70 + 6

OpenStudy (anonymous):

you have to distribute 46/9 to both h and 6